please i really need the answers for tomorrow thank you
well quwstion number one is: find the circumference of a circle that has a radius of 3.2 cm
number two is: find the circumference of a circle that has a radius of 9 cm
number three is: find the circumference of a circle that has a radius of 11.4 cm



Answer by stanbon(75887)

quwstion number one is: find the circumference of a circle that has a radius of 3.2 cm
Ans: C = 2*pi*r
C = 2*3.14159*3.2 = 20.10619.. cm
---------------------------------------------
number two is: find the circumference of a circle that has a radius of 9 cm
C = 2*pi*9 = 56.5487... cm
-----------------------------------------------
number three is:find the circumference of a circle that has a radius of 11.4 cm
C = 2*pi*11.4 = 71.6283.. cm
--------------------------------
Cheers,
Stan H.